OpenStudy (anonymous):

Solve this Separable Equation: x(dy/dx)=1/y^3 with explanations for each step please

OpenStudy (anonymous):

x(dy/dx) = 1/y^3 - for example x * dy = dx / y^3 - communitive property y^3 * x dy = dx - these aren't right y^3 dy = dx/x - but explain each one y^4/4 = ln(x) + C - my teacher is really picky y^4 = 4*ln(x) + 4C - and i dont remember all the terms and lingos y^4 = 4*[ln(x) + C] - integration by parts??? please help

OpenStudy (anonymous):

What's the difficulty here? You have all the steps for a straightforward problem. Is there a specific step you don't follow?

OpenStudy (anonymous):

so for my final answer i got y=+/- (4ln|x|+c)^(1/4) but how do i check if it's correct

OpenStudy (mathmale):

Jamal: I've successfully attached an image of my handwritten solution. Personally, I think you should move on to other problems so that you have more variety in your experience and problem solving skills.

OpenStudy (anonymous):

okay, thank you very much for your help!

OpenStudy (mathmale):

Jamal: Pls go back to our previous discussion; that's where I've attached that image (not here). Very happy to work with you; hope to "see" you again on OpenStudy. Keep moving, gain more experience, save up questions for me (or others on OpenStudy) as you go along.

OpenStudy (mathmale):

Jamal: I've done the whole problem, plus checked the answer, and am attaching an image of that one page. Hope this is helpful. Again, please ask questions or ask for clarification of steps if you need to. All the best to you. MM